# 2cm2cm1cm1cm

### v1

@@ 2.1 Background Lattice
As explained in [12, 13, 14], charge transport in cuprates a priori occurs within two-dimensional isotropic $\mathrm{CuO}_{2}$ layers made of $\mathrm{Cu}^{++}$ and $\mathrm{O}^{--}$ ions, even if a weak inter-layer interaction might also be relevant. The (two-dimensional) lattice composed by copper and oxygen ions is supposed to have the symmetries of the square, as is usual. In other words, it is invariant under the group generated by $90^{\circ}$-degree rotations. This is experimentally verified in the important family of tetragonal cuprates like $\mathrm{La}_{2-x}\mathrm{Sr}_{x}\mathrm{CuO}_{4}$ (LaSr 214) and $\mathrm{La}_{2-x}\mathrm{Ba}_{x}\mathrm{CuO}_{4}$. See, e.g., [25, Section 9.1.2], [12, Section 2.3] and [14, Section 6.3.1].
@@ 2.2 Composite of two Fermions and one Boson
We consider a system of two fermions (electrons or holes in cuprates) with opposite spins interacting via the exchange of one boson in a two dimensional square lattice. Physically, the boson that we have in mind in cuprate superconductors is a spinless bipolaron, since the very strong Jahn-Teller (JT) effect associated with copper ions is an important property of such cuprates [16, 15]. See also [17, 18, 19]. However, the exchanged spinless boson could be of any type, like a phonon or a spin wave, depending on the physical system and mechanism one has in mind.
@@
associated with $\ell^{2}(\mathbb{Z}^{2})$.
We study here the effect of processes of annihilation of two fermions of opposite spins to create a boson, which can conversely be annihilated to create two new fermions. The Hilbert space associated with this composite system, made of two fermions and one boson, is the direct sum $\mathfrak{h}_{f}\oplus\ell^{2}(\mathbb{Z}^{2})$, and not to the tensor product $\mathfrak{h}_{f}\otimes\ell^{2}(\mathbb{Z}^{2})$. This fact already unveils the strong interdependence of the bosonic and fermionic parts. For this reason, from now on we rather use the term “composite of two fermions and one boson” instead of “three-body system”, in order to avoid any misinterpretation.
The fermionic part of the (infinite volume) Hamiltonian of the composite is defined to be the restriction $H_{f}\in\mathcal{B}(\mathfrak{h}_{f})$ of the formal expression
@@
The above fermion-fermion interactions have been extensively studied in condensed matter physics during the last decade, in particular for two dimensional systems. For non-zero functions $\mathrm{u}$, they are named extended Hubbard interactions and they can drastically change the behavior of the system, as compared to the zero-range case (usual Hubbard interaction, $\mathrm{u}=0$). As one example, they are used in the context of ultracold atoms, ions, and molecules [11]. Its bosonic version is also experimentally investigated. See, e.g., [26] published in 2022. In theoretical studies, frequently, only nearest-neighbor interactions added to the on-site (zero-range) Hubbard interactions are considered. Here, we do not need the restriction to one-range (nearest-neighbor) interactions. We only assume that $\mathrm{u}$ is absolutely summable (see (7)), which is physically a very mild restriction, since the effective two-particle repulsive electrostatic potential in crystals is expected to decay exponentially fast in space, because of screening effects.
@@
Similar to the fermionic part, the bosonic part of the (infinite volume) Hamiltonian of the system is defined to be the restriction $H_{b}\in\mathcal{B}(\ell^{2}(\mathbb{Z}^{2}))$ to the one-boson Hilbert space $\ell^{2}(\mathbb{Z}^{2})$ of the formal expression
@@
The term of the Hamiltonian that encodes the decay of a boson into two fermions (i.e., one of the two-electron(hole)-bipolaron-exchange interaction of the Hamiltonian) refers to the bounded operator
@@
$c_{y}^{\ast}\doteq{\sum\limits_{z\in\mathbb{Z}^{2}}}\,\mathfrak{p}\left(z\right)\left(a_{y+z,\uparrow}^{\ast}\,a_{y,\downarrow}^{\ast}+a_{y+z,\uparrow}^{\ast}\,a_{y-z,\downarrow}^{\ast}\right)$ (11)
for some fixed non-zero function $\mathfrak{p}:\mathbb{Z}^{2}\rightarrow\mathbb{R}$ that is invariant under $90^{\circ}$-rotations and exponentially decays in space, i.e.,
${\sum_{z\in\mathbb{Z}^{2}}}\,\mathrm{e}^{\alpha_{0}\left|z\right|}\left|\mathfrak{p}\left(z\right)\right|<\infty\qquad\text{and}\qquad\mathfrak{p}\left(x,y\right)=\mathfrak{p}\left(-y,x\right),\qquad x,y\in\mathbb{Z}\ ,$ (12)
for some $\alpha_{0}>0$. In particular the function $\mathfrak{p}$ is absolutely summable in space. For instance, in [22, 21], $\mathfrak{p}(z)=1$ for $|z|\leq 1$ and $\mathfrak{p}(z)=0$ otherwise. This is the typical example we have in mind, the point here being the fact two fermions on different lattice sites can interact by exchanging a boson. Physically, $c_{y}^{\ast}$ represents the creation of a fermion pair of zero total spin, the two components of which are slightly spread around the lattice position $y\in\mathbb{Z}^{2}$. Such pairs have finite size, because of (12). In fact, as $\mathfrak{p}\neq 0$,
$r_{\mathfrak{p}}\doteq\frac{{\sum_{z\in\mathbb{Z}^{2}}}\,\left|z\right|\left|\mathfrak{p}\left(z\right)\right|}{{\sum_{z\in\mathbb{Z}^{2}}}\,\left|\mathfrak{p}\left(z\right)\right|}\leq\inf_{\alpha_{0}>0}\alpha_{0}^{-1}\sqrt{\frac{{\sum_{z\in\mathbb{Z}^{2}}}\,\mathrm{e}^{\alpha_{0}\left|z\right|}\left|\mathfrak{p}\left(z\right)\right|}{{\sum_{z\in\mathbb{Z}^{2}}}\,\left|\mathfrak{p}\left(z\right)\right|}}<\infty\text{ }$ (13)
@@
Allowing two fermions in different lattice site to interact by exchanging a boson simply means that $r_{\mathfrak{p}}>0$. For the physical significance of this property for cuprates, see [22, 21]. Recall that the exchange strength function $\upsilon:\mathbb{Z}^{2}\rightarrow\mathbb{R}$ is only absolutely summable, not necessarily exponentially decaying as $\mathfrak{p}$, and invariant under $90^{\circ}$-rotations, i.e.,
@@
Note that the Fourier transforms $\hat{\upsilon}$ and $\mathfrak{\hat{p}}$ of $\upsilon$ and $\mathfrak{p}$ are real-valued continuous functions (on the two-dimensional torus $\mathbb{T}^{2}$) that are again invariant under $90^{\circ}$-rotations. Additionally, $\mathfrak{\hat{p}}$ is real analytic, for $\mathfrak{p}$ is exponentially decaying. The reverse process, that is, the annihilation of two unbound fermions to form a boson, is represented by the adjoint operator
@@
The infinite sums (6), (8) and (10) defining formally $H_{f}\in\mathcal{B}\left(\mathfrak{h}_{f}\right)$, $H_{b}\in\mathcal{B}(\ell^{2}(\mathbb{Z}^{2}))$ and $W_{\mathrm{b\rightarrow f}}\in\mathcal{B}(\ell^{2}(\mathbb{Z}^{2}),\mathfrak{h}_{f})$ (9) are to be understood as follows: If $\psi\in\mathfrak{h}_{f}$ or $\psi\in\ell^{2}(\mathbb{Z}^{2})$ is a finitely supported function, then the sum corresponding to $H_{f}\psi$, $H_{b}\psi$ or $W_{\mathrm{b\rightarrow f}}\psi$ is absolutely convergent. Thus, $H_{f}$, $H_{b}$ and $W_{\mathrm{b\rightarrow f}}$ are well-defined linear operators acting on the dense subspace of such functions. One checks that $H_{f}$, $H_{b}$ and $W_{\mathrm{b\rightarrow f}}$ are all bounded on this subspace and they thus have a unique bounded linear extension to the whole Hilbert space where they are defined, namely $\mathfrak{h}_{f}$ for $H_{f}$, and $\ell^{2}(\mathbb{Z}^{2})$ for $H_{b}$ and $W_{\mathrm{b\rightarrow f}}$. We denote the extensions again by $H_{f}$, $H_{b}$ and $W_{\mathrm{b\rightarrow f}}$. Note that, being a bounded operator, $W_{\mathrm{b\rightarrow f}}$ has an adjoint (15), while $H_{f}$ and $H_{b}$ are clearly symmetric and so, self-adjoint, for they are also bounded.
Finally, the full Hamiltonian for the fermion-boson composite is defined, in matrix notation for the direct sum $\mathfrak{h}_{f}\oplus\ell^{2}(\mathbb{Z}^{2})$, as follows:
@@
to this particular subspace $\mathfrak{H}\subseteq\mathfrak{h}_{f}\oplus\ell^{2}(\mathbb{Z}^{2})$. In fact, as the boson is assumed to be spinless, by the conservation of angular momentum, we have that the total spin of the fermion pair resulting from a bosonic decay must be zero. In other words, the physically relevant (vector) states of the fermion-boson compound system always lie in $\mathfrak{H}$. Note finally that $H$ inherits the symmetries of the Hamiltonian (16), i.e., $H$ is invariant under translations and $90^{\circ}$-rotations. Note that this last symmetry, i.e., the rotation invariance, is mainly relevant for the study of unconventional pairings, which is not done here.
@@ 2.3 The Model in Spaces of Quasi-Momenta
We have a composite of two fermions and one boson whose Hamiltonian is translation invariant. In this case, it is a standard procedure (see, e.g. [32, Chapter XIII.16 ]) to use the direct integral decomposition of the Hamiltonian in Fourier space in order to study its spectral properties.
@@
where $|k-p-q|$ is the Euclidean distance between $k$ and $p+q$ in $\mathbb{R}^{2}$. This defines a compact metric space $(\mathbb{T}^{2},d_{\mathbb{T}^{2}})$. Observe also that the usual group operation in $\mathbb{T}^{2}$, i.e., the sum in $\mathbb{R}^{2}$ modulo $(2\pi,2\pi)$, is a continuous operation, while any Borel set in $\mathbb{T}^{2}$ is also a Borel set in $\mathbb{R}^{2}$ (endowed with the Euclidean metric). We also need the normalized Haar measure $\nu$ on $\mathbb{T}^{2}$ defined for any Borel set $B\subseteq\mathbb{T}^{2}$ by
@@
$U:\mathfrak{H}\longrightarrow L^{2}\left(\mathbb{T}^{2},\mathcal{H}\right)$ (24)
in such a way that $k\in\mathbb{T}^{2}$, the fiber quasi-momentum, is exactly the total quasi-momentum of the fermion pair. Recall that $\mathfrak{H}$ (18) is Hilbert space on which $H$ is originally defined. More precisely,
$U\doteq U_{f}\oplus\mathcal{F}\ ,$
@@
$\mathcal{F}:\ell^{2}\left(\mathbb{Z}^{2}\right)\rightarrow L^{2}\left(\mathbb{T}^{2}\right)$
@@
is the composition of two unitary (linear) transformations $U_{1}$ and $U_{2}$, whose exact definitions are given in Section 5.1.
Here, the Fourier transform is defined for any $f\in\ell^{2}(\mathbb{Z}^{2})$ by
@@
$k\cdot x$ being the usual scalar product of $k\in\mathbb{T}^{2}$ and $x\in\mathbb{Z}^{2}$, seen as vectors of $\mathbb{R}^{2}$. We often use the symbol $\widehat{\left(\cdot\right)}$ to shorten the notation of the Fourier transform. For instance, for any $x\in\mathbb{Z}^{2}$, we write $\hat{\mathfrak{e}}_{x}$ to denote the function $\mathrm{e}^{i(\cdot)\cdot x}$ on the torus $\mathbb{T}^{2}$. That is, $\{\hat{\mathfrak{e}}_{x}\}_{x\in\mathbb{Z}^{2}}$ is the image under the Fourier transform of the canonical orthonormal basis$\{\mathfrak{e}_{x}\}_{x\in\mathbb{Z}^{2}}$ (5) of $\ell^{2}(\mathbb{Z}^{2})$.

### v2

@@ 2.1 Background Lattice
Copper oxide superconductors have a relatively complex three-dimensional lattice structure. However, they always contain parallel two-dimensional layers of copper ($\mathrm{Cu}^{++}$) and oxygen ($\mathrm{O}^{--}$) ions. These $\mathrm{CuO}_{2}$ layers are essential to understanding low-temperature superconducting properties, because the (superconducting) charge transport takes place within the layers. This is explained in [12, 13, 14]. Considering a weak inter-layer interaction might also help to increase prediction accuracy, but charge transport between each $\mathrm{CuO}_{2}$ layer or, more generally, in the direction orthogonal to each layer remains negligible [^4] The superconducting coherence length is much smaller in this orthogonal direction than in the parallel planes made of copper and oxygen ions.. Each $\mathrm{CuO}_{2}$ layer generally has the symmetries of the square. In other words, it is invariant under the group $\{0,\pi/2,\pi,3\pi/2\}$ generated by $90^{\circ}$-degree rotations. See, e.g., [31, Section 9.1.2], [12, Section 2.3] and [14, Section 6.3.1]. This is an important symmetry property that we keep in mind throughout our study.
@@ 2.2 Composite of two Fermions and one Boson
We consider a system of two fermions (electrons or holes in cuprates) with opposite spins interacting via the exchange of one boson in a two dimensional square lattice. Physically, the boson that we have in mind in cuprate superconductors is a spinless bipolaron, since the very strong Jahn-Teller (JT) effect associated with copper ions is an important property of such cuprates [16, 15]. See Section 5.1 for more details. However, the exchanged spinless boson could be of any type, like a phonon or a spin wave, depending on the physical system and mechanism one has in mind.
[…]
@@
associated with $\ell^{2}(\mathbb{Z}^{2})$. See Equation (155) below for the precise definition of $\mathfrak{F}_{+}$. For a concise review of bosonic and fermionic Fock spaces, as well as the corresponding annihilation and creation operators, see Section 5.2.
We study here the effect of processes of annihilation of two fermions of opposite spins to create a boson, which can conversely be annihilated to create two new fermions. The Hilbert space associated with this composite system, made of two fermions and one boson, is the direct sum $\mathfrak{h}_{f}\oplus\ell^{2}(\mathbb{Z}^{2})$, and not the tensor product $\mathfrak{h}_{f}\otimes\ell^{2}(\mathbb{Z}^{2})$. Note indeed that $\mathfrak{h}_{f}\oplus\ell^{2}(\mathbb{Z}^{2})$ can naturally be identified [^8] Denote the vacuum of the Fock space $\mathfrak{F}_{\pm}$ by $\Omega_{\pm}$ and define the mapping $\varsigma$ from $\mathfrak{h}_{f}\oplus\ell^{2}(\mathbb{Z}^{2})$ to $\mathfrak{F}_{-}\otimes\mathfrak{F}_{+}$ by $\varsigma\left(\varphi\oplus\psi\right)=\varphi\otimes\Omega_{-}+\Omega_{+}\otimes\psi$. Then, observe that $\varsigma$ is an isometric linear transformation from $\mathfrak{h}_{f}\oplus\ell^{2}(\mathbb{Z}^{2})$ to $\mathfrak{F}_{-}\otimes\mathfrak{F}$. with a subspace of $\mathfrak{F}_{-}\otimes\mathfrak{F}_{+}$. This fact already unveils the strong interdependence of the bosonic and fermionic parts. For this reason, from now on we rather use the term “composite of two fermions and one boson” instead of “three-body system”, in order to avoid any misinterpretation.
Fermionic Hamiltonian. The fermionic part of the (infinite volume) Hamiltonian of the composite is defined to be the restriction $H_{f}\in\mathcal{B}(\mathfrak{h}_{f})$ of the formal expression
@@
Extended Hubbard interactions. The above fermion-fermion interactions have been extensively studied in condensed matter physics during the last decade, in particular for two dimensional systems. For non-zero functions $\mathrm{u}$, they are named extended Hubbard interactions and they can drastically change the behavior of the system, as compared to the zero-range case (usual Hubbard interaction, $\mathrm{u}=0$). As one example, they are used in the context of ultracold atoms, ions, and molecules [11]. Its bosonic version is also experimentally investigated. See, e.g., [32] published in 2022.
In theoretical studies, frequently, only nearest-neighbor interactions added to the on-site (zero-range) Hubbard interactions are considered. Here, we do not need the restriction to one-range (nearest-neighbor) interactions. We only assume that $\mathrm{u}$ is absolutely summable (see (7)), which is physically a very mild restriction, since the effective two-particle repulsive electrostatic potential in crystals is expected to decay exponentially fast in space, because of screening effects.
@@
Bosonic Hamiltonian. Similar to the fermionic part, the bosonic part of the (infinite volume) Hamiltonian of the system is defined to be the restriction $H_{b}\in\mathcal{B}(\ell^{2}(\mathbb{Z}^{2}))$ to the one-boson Hilbert space $\ell^{2}(\mathbb{Z}^{2})$ of the formal expression
@@
Exchange interactions. The term of the Hamiltonian that encodes the decay of a boson into two fermions (i.e., one of the two-electron(hole)-bipolaron-exchange interaction of the Hamiltonian) refers to the bounded operator
@@
$c_{y}^{\ast}\doteq{\sum\limits_{z\in\mathbb{Z}^{2}}}\,\left(\mathfrak{p}_{1}\left(z\right)a_{y+z,\uparrow}^{\ast}\,a_{y,\downarrow}^{\ast}+\mathfrak{p}_{2}\left(2z\right)a_{y+z,\uparrow}^{\ast}\,a_{y-z,\downarrow}^{\ast}\right)$ (11)
for some fixed functions $\mathfrak{p}_{1},\mathfrak{p}_{2}:\mathbb{Z}^{2}\rightarrow\mathbb{R}$ that are invariant under $90^{\circ}$-rotations and exponentially decay in space, i.e.,
${\sum_{z\in\mathbb{Z}^{2}}}\,\mathrm{e}^{\alpha_{0}\left|z\right|}\left|\mathfrak{p}_{\sharp}\left(z\right)\right|<\infty\qquad\text{and}\qquad\mathfrak{p}_{\sharp}\left(x,y\right)=\mathfrak{p}_{\sharp}\left(-y,x\right),\qquad x,y\in\mathbb{Z},\ \sharp\in\{1,2\}\ ,$ (12)
for some $\alpha_{0}>0$. In particular, the functions $\mathfrak{p}_{1},\mathfrak{p}_{2}$ are absolutely summable in space.
By definition, we take $\mathfrak{p}_{2}\left(z\right)\doteq 0$ if $z\in\mathbb{Z}^{2}\backslash(2\mathbb{Z})^{2}$ and we also assume that
$\mathfrak{p}_{1}+\mathfrak{p}_{2}\neq 0\text{\qquad}\mathrm{and}\text{\qquad}\mathfrak{p}_{2}\left(x\right)\neq-\mathrm{e}^{i\frac{k}{2}\cdot x}\mathfrak{p}_{1}\left(x\right)\ ,\qquad x\in\mathbb{Z},\ k\in[-\pi,\pi)^{2}\ .$ (13)
The condition $\mathfrak{p}_{1}+\mathfrak{p}_{2}\neq 0$ only ensures the non-triviality of the exchange interaction, while the second condition avoids the singular case of a quasi-momentum $k_{0}\in[-\pi,\pi)^{2}$ at which the exchange interaction trivially vanishes, see below (36). This case can easily be analyzed, but it makes the argumentation cumbersome, so we omit it here, as it is a highly unusual and irrelevant situation. For example, (13) is already satisfied as soon as $\mathfrak{p}_{1}\left(z\right)\neq 0$ for some $z\in\mathbb{Z}^{2}\backslash(2\mathbb{Z})^{2}$, since $\mathfrak{p}_{2}\left(z\right)\doteq 0$ for any $z\notin(2\mathbb{Z})^{2}$. In [22, Eq. (6)], $\mathfrak{p}_{2}=0$ and, given $\kappa>0$, $\mathfrak{p}_{1}(z)=\mathrm{e}^{-\kappa|z|}$ for $|z|\leq 1$ and $\mathfrak{p}_{1}(z)=0$ otherwise, while in [21, Eq. (4)], $\mathfrak{p}_{2}\left(2z\right)=\mathfrak{p}_{1}\left(z\right)=1$ when $|z|\leq 1$ and $\mathfrak{p}_{1}\left(z\right)=\mathfrak{p}_{2}\left(z\right)=0$ otherwise. This are the typical examples we have in mind, the point here being the fact two fermions on different lattice sites can interact by exchanging a boson. See also Section 5.1.
Physically, $c_{y}^{\ast}$ represents the creation of a fermion pair of zero total spin, the two components of which are slightly spread around the lattice position $y\in\mathbb{Z}^{2}$. Such pairs have finite size, because of (12). In fact,
$r_{\mathfrak{p}}\doteq\frac{1}{2}\left(r_{\mathfrak{p}_{1}}+r_{\mathfrak{p}_{2}}\right)\ ,$ (14)
where, for any $\sharp\in\{1,2\}$, $r_{\mathfrak{p}_{\sharp}}\doteq 0$ if $\mathfrak{p}_{\sharp}=0$, otherwise it is equal to
$r_{\mathfrak{p}_{\sharp}}\doteq\frac{{\sum_{z\in\mathbb{Z}^{2}}}\,\left|z\right|\left|\mathfrak{p}_{\sharp}\left(z\right)\right|}{{\sum_{z\in\mathbb{Z}^{2}}}\,\left|\mathfrak{p}_{\sharp}\left(z\right)\right|}\leq\inf_{\alpha_{0}>0}\alpha_{0}^{-1}\sqrt{\frac{{\sum_{z\in\mathbb{Z}^{2}}}\,\mathrm{e}^{\alpha_{0}\left|z\right|}\left|\mathfrak{p}_{\sharp}\left(z\right)\right|}{{\sum_{z\in\mathbb{Z}^{2}}}\,\left|\mathfrak{p}_{\sharp}\left(z\right)\right|}}<\infty\ ,$ (15)
@@
Allowing two fermions in different lattice site to interact by exchanging a boson simply means that $r_{\mathfrak{p}}>0$. For the physical significance of this property for cuprates, see [22, 21].
Recall that the exchange strength function $\upsilon:\mathbb{Z}^{2}\rightarrow\mathbb{R}$ is only absolutely summable, not necessarily exponentially decaying as $\mathfrak{p}_{1}$ and $\mathfrak{p}_{2}$, and invariant under $90^{\circ}$-rotations, i.e.,
@@
Note that the Fourier transforms $\hat{\upsilon}$,$\mathfrak{\hat{p}}_{1}$ and $\mathfrak{\hat{p}}_{2}$ of $\upsilon$, $\mathfrak{p}_{1}$ and $\mathfrak{p}_{2}$ are real-valued continuous functions (on the two-dimensional torus $\mathbb{T}^{2}$) that are again invariant under $90^{\circ}$-rotations. Additionally, $\mathfrak{\hat{p}}_{1}$ and $\mathfrak{\hat{p}}_{2}$ are real analytic, for $\mathfrak{p}_{1}$ and $\mathfrak{p}_{2}$ are exponentially decaying. The reverse process, that is, the annihilation of two unbound fermions to form a boson, is represented by the adjoint operator
@@
Mathematical remarks. The infinite sums (6), (8) and (10) defining formally $H_{f}\in\mathcal{B}\left(\mathfrak{h}_{f}\right)$, $H_{b}\in\mathcal{B}(\ell^{2}(\mathbb{Z}^{2}))$ and $W_{\mathrm{b\rightarrow f}}\in\mathcal{B}(\ell^{2}(\mathbb{Z}^{2}),\mathfrak{h}_{f})$ (9) are to be understood as follows: If $\psi\in\mathfrak{h}_{f}$ or $\psi\in\ell^{2}(\mathbb{Z}^{2})$ is a finitely supported function, then the sum corresponding to $H_{f}\psi$, $H_{b}\psi$ or $W_{\mathrm{b\rightarrow f}}\psi$ is absolutely convergent. Thus, $H_{f}$, $H_{b}$ and $W_{\mathrm{b\rightarrow f}}$ are well-defined linear operators acting on the dense subspace of such functions. One checks that $H_{f}$, $H_{b}$ and $W_{\mathrm{b\rightarrow f}}$ are all bounded on this subspace and they thus have a unique bounded linear extension to the whole Hilbert space where they are defined, namely $\mathfrak{h}_{f}$ for $H_{f}$, and $\ell^{2}(\mathbb{Z}^{2})$ for $H_{b}$ and $W_{\mathrm{b\rightarrow f}}$. We denote the extensions again by $H_{f}$, $H_{b}$ and $W_{\mathrm{b\rightarrow f}}$. Note that, being a bounded operator, $W_{\mathrm{b\rightarrow f}}$ has an adjoint (17), while $H_{f}$ and $H_{b}$ are clearly symmetric and so, self-adjoint, for they are also bounded.
Full model. Finally, the full Hamiltonian for the fermion-boson composite is defined, in matrix notation for the direct sum $\mathfrak{h}_{f}\oplus\ell^{2}(\mathbb{Z}^{2})$, as follows:
@@
to this particular subspace $\mathfrak{H}\subseteq\mathfrak{h}_{f}\oplus\ell^{2}(\mathbb{Z}^{2})$.
In fact, as the boson is assumed to be spinless, by the conservation of angular momentum, we have that the total spin of the fermion pair resulting from a bosonic decay must be zero. In other words, the physically relevant (vector) states of the fermion-boson compound system always lie in $\mathfrak{H}$. Note finally that $H$ inherits the symmetries of the Hamiltonian (18), i.e., $H$ is invariant under translations and $90^{\circ}$-rotations. Note that this last symmetry, i.e., the rotation invariance, is mainly relevant for the study of unconventional pairings, which is not done here.
@@ 2.3 The Model in Spaces of Quasi-Momenta
We have a composite of two fermions and one boson whose Hamiltonian is translation invariant. In this case, it is a standard procedure (see, e.g. [38, Chapter XIII.16 ]) to use the direct integral decomposition of the Hamiltonian in Fourier space in order to study its spectral properties.
@@
where $|k-p-q|$ is the Euclidean distance between $k$ and $p+q$ in $\mathbb{R}^{2}$. This defines a compact metric space $(\mathbb{T}^{2},d_{\mathbb{T}^{2}})$. Observe also that the usual group operation in $\mathbb{T}^{2}$, i.e., the sum in $\mathbb{R}^{2}$ modulo $(2\pi,2\pi)$, is a continuous operation, while any Borel set in $\mathbb{T}^{2}$ is also a Borel set in $\mathbb{R}^{2}$ (endowed with the Euclidean metric).
We also need the normalized Haar measure $\nu$ on $\mathbb{T}^{2}$ defined for any Borel set $B\subseteq\mathbb{T}^{2}$ by
@@
$\mathbb{U}:\mathfrak{H}\longrightarrow L^{2}\left(\mathbb{T}^{2},\mathcal{H}\right)$
in such a way that $k\in\mathbb{T}^{2}$, the fiber quasi-momentum, is exactly the total quasi-momentum of the fermion pair.
Recall that $\mathfrak{H}$ defined in (20) is the Hilbert space on which $H$ is originally defined. More precisely,
$\mathbb{U}\doteq U_{f}\oplus\mathcal{F}\ ,$ (26)
@@
$\mathcal{F}:\ell^{2}\left(\mathbb{Z}^{2}\right)\rightarrow L^{2}\left(\mathbb{T}^{2}\right)$ (27)
@@
is the composition of two unitary (linear) transformations $U_{1}$ and $U_{2}$, whose exact definitions are given as follows:
$\begin{array}[]{cccccccc}U_{1}:&\mathfrak{h}_{0}&\rightarrow&\ell^{2}\left(\mathbb{Z}^{2}\times\mathbb{Z}^{2}\right)&\rightarrow&\ell^{2}\left(\mathbb{Z}^{2}\times\mathbb{Z}^{2}\right)&\rightarrow&\ell^{2}\left(\mathbb{Z}^{2}\right)\otimes\ell^{2}\left(\mathbb{Z}^{2}\right)\\ &\mathfrak{e}_{(x,\uparrow)}\wedge\mathfrak{e}_{(y,\downarrow)}&\mapsto&\mathfrak{e}_{(x,y)}&\mapsto&\mathfrak{e}_{(x,x-y)}&\mapsto&\mathfrak{e}_{x}\otimes\mathfrak{e}_{x-y}\end{array}$ (29)
and
$\begin{array}[]{cccccc}U_{2}:&\ell^{2}\left(\mathbb{Z}^{2}\right)\otimes\ell^{2}\left(\mathbb{Z}^{2}\right)&\rightarrow&L^{2}\left(\mathbb{T}^{2}\right)\otimes L^{2}\left(\mathbb{T}^{2}\right)&\rightarrow&\int_{\mathbb{T}^{2}}^{\oplus}L^{2}\left(\mathbb{T}^{2}\right)\nu\left(\mathrm{d}k\right)\\ &\mathfrak{e}_{x}\otimes\mathfrak{e}_{x-y}&\mapsto&\mathfrak{\hat{e}}_{x}\otimes\mathfrak{\hat{e}}_{x-y}&\mapsto&\mathfrak{\hat{e}}_{x}\left(\cdot\right)\mathfrak{\hat{e}}_{x-y}\end{array}\ .$ (30)
Because $\{\mathfrak{e}_{(x,\uparrow)}\wedge\mathfrak{e}_{(y,\downarrow)}\}_{x,y\in\mathbb{Z}^{2}}$, $\{\mathfrak{e}_{(x,y)}\}_{x,y\in\mathbb{Z}^{2}}$ and $\{\mathfrak{e}_{x}\otimes\mathfrak{e}_{y}\}_{x,y\in\mathbb{Z}^{2}}$ are orthonormal bases and $(x,y)\mapsto(x,x-y)$ is a bijection on $\mathbb{Z}^{2}\times\mathbb{Z}^{2}$, $U_{1}$ is well-defined as a composition of three unitary linear transformations. Note also that the last unitary linear transformation defining $U_{2}$ is defined as in Proposition 5.8, while the first one defining $U_{2}$ is the tensor product $\mathcal{F}\otimes\mathcal{F}$ of the Fourier transform $\mathcal{F}$ on $\ell^{2}(\mathbb{Z}^{2})$, defined for any $f\in\ell^{1}(\mathbb{Z}^{2})\subseteq\ell^{2}(\mathbb{Z}^{2})$ by
@@
$k\cdot x$ being the usual scalar product of $k\in\mathbb{T}^{2}$ and $x\in\mathbb{Z}^{2}$, seen as vectors of $\mathbb{R}^{2}$. Here we use the symbol $\widehat{\left(\cdot\right)}$ to shorten the notation of the Fourier transform. For instance, for any $x\in\mathbb{Z}^{2}$, we write above $\hat{\mathfrak{e}}_{x}$ to denote the function $\mathrm{e}^{i(\cdot)\cdot x}$ on the torus $\mathbb{T}^{2}$. That is, $\{\hat{\mathfrak{e}}_{x}\}_{x\in\mathbb{Z}^{2}}$ is the image under the Fourier transform of the canonical orthonormal basis$\{\mathfrak{e}_{x}\}_{x\in\mathbb{Z}^{2}}$ (5) of $\ell^{2}(\mathbb{Z}^{2})$.